OpenStudy (anonymous):

Increase 62 by 40% Decrease 89 by 60%

OpenStudy (anonymous):

62*40/100 = 24.8 , ==> 62+24.8=86.8 ===== 89*60/100 = 53.4 ===> 89-53.4=25.6